OpenStudy (anonymous):

sqrt of 96 8 sqrt 63x^5 sqrt128x^5y^2 ^3 sqrt 32 ^3 sqet 56x^14

OpenStudy (anonymous):

how do you get these not in decimal form

OpenStudy (anonymous):

you must find a factor of 96 which is a perfect square

OpenStudy (anonymous):

i dont know how to do that without a calculator?

OpenStudy (anonymous):

\[\sqrt{96}= \sqrt{(4)(4)(6)}=4\sqrt{6}\]